Find: 15% of 250 Find: 15% of 250 Find: 15% of 250 Find: 15% of 250

 Find: 
15% of 250
Find: 
15% of 250
Find: 
15% of 250
Find: 
15% of 250

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (15/100) × 250
= (15/10) × 25
= (15/2) × 5
= (75/2)
= 37.5

Thanks